# PANTS: Practical Adversarial Network Traffic Samples against ML-powered Networking Classifiers

### v1

@@ 5.2 Key Findings
Importantly, PANTS’s ASR is more consistent across runs in all the test cases. This is important as an operator is unlikely to trust the assessment of a tool that reports a different ASR in every run i.e., varying levels of vulnerability for a given MNC. Concretely, the average standard deviation across all the test cases is 0.77 for PANTS, while for Amoeba is 3.12 and BAP is 2.61. For some specific test cases, Amoeba and BAP show ASR with relatively large fluctuations. For example, Amoeba’s ASR on attacking TF ranges from 76.15% to 93.45% and BAP is from 43.89% to 54.25% while PANTS’ ASR only ranges from 99.75% to 100.0% as we observe in Fig. 6(a). The large variance in Amoeba’s results stems from a well-known problem in RL: its inherent instability of training. For PANTS, randomness only comes from the adversarial features generated by PGD and its timeout mechanism. PGD might generate different adversarial features in different runs due to random restarts, which are used in MLP and RF [27]. The time budget in SMT also causes minor randomness by the CPU’s processing speed fluctuation. It turns out that randomness caused by these two factors are small.
@@
Finding 2: Iterative augmentation with adversarial, realizable and semantics-preserving samples improves the robustness of an MNC without hurting its accuracy.
As Amoeba is more general and results in higher ASR compared to BAP, we continue our evaluation on robustification with Amoeba only. Concretely, we robustify the target ML models by iteratively fine-tuning them on a combination of the original dataset with its adversarial counterpart generated by PANTS and Amoeba as the way described in §4.5.
We found that in most of the cases, 3 iterations of robustification are enough when using PANTS to robustify an MNC. The robustified MNC is able to defend the adversarial inputs that are generated by PANTS itself, Amoeba and BAP. For TF in the application APP and QOE under end-host threat model, we train for 5 iterations to reach a desirable ASR. Figure 7 shows a trajectory of ASR when a vanilla MNC is in iterative robustification using PANTS’ samples. During the iterative training the ASR (calculated using PANTS to generate adversarial inputs) drops, meaning that the model is getting more robust. Iteratively training MNCs using adversarial samples generated by Amoeba follows a similar but less dramatic trend. The results are shown in Appendix in the fig. 12.
(a) Robustified for defending end-host attackers.
(b) Robustified for defending in-path attackers.
Figure 7: ASR calculated by PANTS while the MNC is trained for the i-th time leveraging a dataset augmented with adversarial samples produced by PANTS. MNCs are getting more robust against adversarial inputs over time.
After robustification, PANTS generates multiple adversarial samples for each benign counterpart in total to cover diverse adversarial samples. Using PANTS, we generate around 1.16-180 samples for each benign sample during robustification based on different threat models, applications and ML models. We also include non-adversarial samples generated by PANTS (not just adversarial). This is beneficial for training because the non-adversarial samples still have a higher loss than their original benign counterpart, being closer to the decision boundary, even though they can be correctly classified. Including those samples can help form a more refined decision boundary. In any case, including non-adversarial samples in the training set offers Amoeba more data to train.
@@
We compare four ways of robustification, ①: authentic adversarial-training, and iterative augmentation using ②: PGD’s samples, ③: PANTS’ samples, ④: Amoeba’s samples. ②, ③ and ④ share the same iterative way of robustification and the only difference is the method to generate adversarial samples. PANTS and Amoeba generate adversarial, semantic-preserving, and adversarial samples while PGD just generates adversarial samples. Figure 8 shows the accuracy and ASR of the robustfied ML models against end-host attackers under the application APP. Adversarially-trained models are robust against attacks from PANTS and Amoeba, but suffer from low accuracy. In contrast, iterative augmentation doesn’t drop the model accuracy, but not all of them can improve robustness for the model. PANTS-robustified models are more robust against both PANTS and Amoeba compared to the vanilla one. But Amoeba-robustified models can only be robust against Amoeba and PGD-robustified cannot be robust against any of the attacks.
Finding 3: PANTS improves the robustness of a model against any attack method 142% more than Amoeba
@@
Figure 9: ASR for the vanilla, PANTS-robustified and Amoeba-robustified models for the end-host threat model. PANTS-robustified’s ASR is the lowest, showing PANTS’s ability to generate adversarial samples that effectively robustify target models.
@@
Table 2: Average robustness improvement. PANTS-robustified models have 142% more robustness improvement on average compared to Amoeba-robustified ones.
@@
After robustification, we use again PANTS, Amoeba and BAP to re-attack (generate adversarial inputs for) the robustified models $M^{\prime}$ to re-evaluate models’ vulnerability. We train a new agent for Amoeba on $D_{train}^{\prime}$ to fool each robustified model $M^{\prime}$. Figure 9 shows the ASR using PANTS, Amoeba, or BAP to find adversarial samples from the vanilla, PANTS, and Amoeba robustified models for the end-host threat model. ASR for the in-path threat model is shown in figure 11 in the Appendix. These results show that PANTS-robustified models are generally robust against the attack from PANTS, Amoeba and BAP, while Amoeba-robustified models are only robust against Amoeba. Table 2 shows the robustness improvement when robustified by PANTS and Amoeba and evaluated with PANTS, Amoeba and BAP. Robustness improvement is defined as $(ASR_{vanilla}-ASR_{robustified})/ASR_{vanilla}\times 100\%$. PANTS-robustified models demonstrate 52.72% robustness improvement on average while Amoeba-robustified ones only demonstrate 21.76%. PANTS-robustified models can have 142% more robustness improvement than Amoeba-robustified ones.
Prior works, especially in RL, have found that adversarial training can only defend a model against adversarial samples generated by the same generation process as the one used in training [37]. However, we found that augmenting with a large number of adversarial, realizable and semantics-preserving samples has the potential to robustify the model against multiple different attacks. That is the case because the space of adversarial, realizable, and semantics-preserving samples (generated by PANTS) is smaller than the space of samples that are only adversarial.
@@
PANTS-robustified models demonstrate considerable robustness improvement against PANTS, Amoeba and BAP. However, we do not claim that PANTS-robustified models are robust against any possible attacks. First, the ASR for PANTS-robustified models doesn’t drop to 0, meaning that these attack methods can still find adversarial samples against the robustified models. Second, as an empirical way of robustification, we don’t have a theoretical guarantee that the robustified models are robust against the full spectrum of perturbation. Figure 11(f) shows that there still exists at least one such that the PANTS-robustified model doesn’t provide enough robustness. Still, this work presents a clear, extendible move towards the right direction.
Finding 4: Transferability provides an opportunity to robustify a MNC without white-box access.
@@
Figure 10: Transferability results for end-host threat model for the application APP using PANTS’ generated adversarial samples. Most of the transfer ASR are higher than 50%, indicating that the adversarial samples generated for one implementation of an MNC is adversarial others.
@@
There are use cases in which a network operator will not have access to the MNC,   e.g., because the MNC is designed by a third party and its details are secret. While this is not our target setting, we show that transferability of adversarial samples PANTS might be able to help the network operator e.g., by providing examples in which the target MNC misclassified or by providing potentially intresting test cases for the MNC designer.
@@
Transferability is defined as the ratio of adversarial samples that are generated for a given implementation of ML classifier that can also be adversarial against another one. Concretely, suppose we evaluate transferability of the adversarial samples generated from $M_{A}$ to $M_{B}$, we first construct the test set $X_{test\_A\_B}^{\prime}\subseteq X_{test}$ that both $M_{A}$ and $M_{B}$ can correctly classify. Then we use PANTS to generate adversarial samples for $X_{test\_A\_B}^{\prime}$ using $M_{A}$ and evaluate its adversariness against $M_{B}$. Figure 10 shows the transferability result between MLP, RF, TF and CNN for end-host threat model tested for the application APP. Most of the transfer ASR is above 50%, showing that PANTS can be useful even without full access to the MNC implementation. To benefit from it the operator would need to train a proxy model and find adversarial samples against it using PANTS. According to Figure 10, the adversarial samples against the proxy model are likely to be adversarial against the target one. Note that the transfer ASR between different models cannot be compared directly. For example, the ASR of transferring adversarial samples from CNN to RF is 70.44% while directly attacking RF (transfer ASR from RF to RF) is 66.5%. This doesn’t necessarily mean that transferring CNN’s adversarial samples to RF is more effective than directly attacking RF directly attacking RF because transferring from CNN to RF is tested on $X_{test\_CNN\_RF}^{\prime}$ while directly attack RF is tested on $X_{test\_RF\_RF}^{\prime}$. As $X_{test\_CNN\_RF}^{\prime}\subset X_{test\_RF\_RF}^{\prime}$, the two tested datasets may not be identical.
@@ 6 Conclusion
Appendix A Features used in different applications
@@
Appendix B Hyperparemeters of ML models
@@
Appendix C Capability extension of Amoeba
@@
Appendix D Testbed specification
For consistent evaluation, all the experiments using PANTS to generate adversarial samples are running on the cluster of Cloudlab machines [12]. Each machine has two Intel Xeon Silver 4114 10-core CPUs at 2.20 GHz and 192GB ECC DDR4-2666 memory. The rest of the experiments related to Amoeba rely on GPU training, which is running on two Azure machines [3]. Each one uses 16 vCPUs from AMD EPYC 7V12(Rome) CPU, 110GB memory, and an Nvidia Tesla T4 GPU with 16 GB GPU memory.
Appendix E Robustness of the PANTS- and Amoeba-robustified models
Figure 11 shows the robustness of using PANTS, Amoeba and BAP to attack PANTS- and Amoeba-robustified models under the in-path threat model.
@@ Appendix E Robustness of the PANTS- and Amoeba-robustified models
Figure 11: ASR for the vanilla, PANTS-robustified and Amoeba-robustified models for the in-path threat model. PANTS-robustified’s ASR is lower, demonstrating PANTS’s ability to generate adversarial samples that effectively robustify target models. For the dataset QOE, using Amoeba for robustification is skipped since it cannot effectively generate enough adversarial samples.

### v3

@@ 5.2 Key Findings
Further, PANTS’s ASR is more consistent across runs in all the test cases. This is important as an operator is unlikely to trust the assessment of a tool that reports a different ASR in every run i.e., varying levels of vulnerability for a given MNC. Concretely, the average standard deviation across all the test cases is 0.77 for PANTS, while for Amoeba 3.12 and for BAP 2.61. For some cases, specifically, Amoeba and BAP show much higher fluctuations in ASR compared to PANTS. For example, Amoeba’s ASR on attacking TF ranges from 76.15% to 93.45% and BAP is from 43.89% to 54.25% while PANTS’ ASR only ranges from 99.75% to 100.0% as we observe in Fig.  6(a). The large variance in Amoeba’s results stems from a well-known problem in RL: its inherent instability of training. BAP also struggles with instability when training the perturbation generator but is better than Amoeba. For PANTS, randomness is limited only stemming from (i) PGD restarts used in MLP and RF [37]; and (ii) CPU fluctuations affecting the SMT solver’s computation speed.
@@
  2 Iterative augmentation with adversarial, realizable, and semantics-preserving samples improves the robustness of an MNC without hurting its accuracy.
To evaluate our iterative adversarial, we compare four approaches for robustifying MNCs. First, we include authentic adversarial training, which integrates PGD into the training process. Next, we include three versions of our iterative augmentation explained in §4.5, where adversarial examples are generated by PGD, PANTS, and Amoeba. Examples generated by PANTS and Amoeba are also semantic-preserving. Fig. 7 shows the accuracy and ASR of the robustfied ML models against end-host attackers under the application APP. Authentic adversarial training enhances the robustness of the models against attacks from PANTS (Fig. 7(a)) and Amoeba Fig. 7(b), but hurts their accuracy. In contrast, iterative augmentation (our approach) maintains the models’ high accuracy (as shown in Table  1) and could improve robustness subject to the quality of the generated samples. Concretely, when samples are generated by PGD, iterative augmentation does not help in robustifying the targeted models. When samples are generated by Amoeba, iterative augmentation robustifies models against samples generated by Amoeba. Finally, when samples are generated by PANTS (i.e., as in the complete PANTS framework), iterative augmentation makes models robust against both PANTS and Amoeba. Fig. 8 and Fig. 16 offer a more detailed view of this result by plotting the improvement of ASR as a MNC is under the iterative robustification process. clearly, ASR decreases more dramatically when samples are generated by PANTS compared to Amoeba.
@@
(a) Robustified for defending end-host attackers.
(b) Robustified for defending in-path attackers.
Figure 8: ASR calculated by PANTS while the MNC is trained for the i-th time leveraging a dataset augmented with adversarial samples produced by PANTS. MNCs are getting more robust against adversarial inputs over time. For the results using Amoeba’s samples for augmentation, please refer to Fig. 16 in the Appendix.
  3 PANTS improves the robustness tested MNCs 142% more than Amoeba and ~40X more than NetShare.
Refer to caption (a) Accuracy of RF
Refer to caption (b) Vulnerability of RF
Refer to caption (c) Accuracy of CNN
Refer to caption (d) Vulnerability of CNN
Figure 9: Accuracy and ASR when training two APP implementations with synthetic data generated by NetShare and adversarial data by PANTS. NetShare cannot improve the robustness of MNCs. Similar results for the other two implementations are reported in Fig. 14 in the Appendix.
@@
Figure 10: ASR for the vanilla, PANTS-robustified and Amoeba-robustified models for the end-host threat model. PANTS-robustified’s ASR is the lowest. Similar results for in-path threat model are reported in Fig. 15 in the Appendix.
@@
Table 2: Average robustness improvement. PANTS-robustified models are 142% more robust on average compared to Amoeba-robustified ones.
@@
After robustification, we use again PANTS, Amoeba and BAP to re-attack (spawn adversarial inputs for) the robustified models $M^{\prime}$ to re-evaluate models’ vulnerability. We train a new agent for Amoeba on $D_{train}^{\prime}$ to fool each robustified model $M^{\prime}$. Fig. 10 shows the ASR using PANTS, Amoeba, or BAP to find adversarial samples from the vanilla, PANTS, and Amoeba robustified models for the end-host threat model (refer to Fig. 15 for the in-path threat model). These results show that PANTS-robustified models are generally robust against the attack from PANTS, Amoeba and BAP, while Amoeba-robustified models are only robust against Amoeba. Table 2 shows the robustness improvement when robustified by PANTS and Amoeba and evaluated with PANTS, Amoeba and BAP. PANTS-robustified models are 52.72% more robust on average than the vanilla ones while Amoeba-robustified ones are only 21.76% more robust. PANTS-robustified models outperform Amoeba-robustified by 142%.
We also compare PANTS’s ability to robustify MNCs with NetShare, which (similar to PANTS) generates synthetic data, albeit non-adversarial. Fig. 9 and Fig. 14 show APP’s accuracy and ASR for end-host threat model implemented with PANTS when using the same amount of data generated by PANTS and NetShare for robustification. PANTS-robustified models have slightly lower accuracy than NetShare-robustified ones. However, the accuracy of PANTS-robustified models is still high, meaning that they are accurate enough against common inputs and they are more robust against adversarial perturbation than NetShare-robustified ones. NetShare can hardly improve the models’ robustness.
  4 PANTS can improve the robustness of MNCs even against threat models outside those used during adversarial training.
Refer to caption Figure 11: The ASR of two APP implementations for various threat models, before and after they were robustified by PANTS against our end-host attacker (20,20,20,0). PANTS improves the robustness of MNCs against threat models that are different or even stronger than what was considered during robustification.
We evaluate PANTS ability to robustify MNCs against a more diverse set of attackers, including attackers that are strictly stronger than what was used during robustification. To this end, we construct $16$ attackers, shown in Fig. 11, which vary on the attacker’s access to the model and their capabilities. The notation w(b) refers to the white-box(black-box) attacker, while the ($\alpha$, $\beta$, $\gamma$, $\delta$) represents an attacker who can delay packets with at most $\alpha\%$, append payload for at most $\beta\%$ packets, inject at most $\gamma$ packets, and split at most $\delta\%$ packets. [^3] For each packet, we only allow to split at most once to prevent a packet split into too many small fragments. Fig. 11 reports the ASR for MLP and CNN for the application APP before and after each model has been robustified by PANTS. Observe that all models are robustified against the end-host attackers described in §5.1 (i.e., a (20, 20, 20, 0) attacker), meaning that many of the attackers tested are stronger. Importantly, PANTS is able to improve the robustness of these models even against strictly stronger attackers. For instance, PANTS-robustified MLP is 72% more robust than the vanilla version against an attacker that can delay, inject and append twice the amount of the attacker used for robustification and can also split packets. This happens because during adversarial training, even assuming a weaker or different adversary, the model is guided away from non-robust features (i.e., features that are not strongly correlated with the output but rather coincidental or dataset-specific) towards meaningful and resilient features which encapsulate genuine semantics [28].
@@
While PANTS-robustified models demonstrate considerable robustness improvement against PANTS, Amoeba and BAP (as discussed in Findings 3 and 4), we do not claim that PANTS-robustified models are robust against any possible attacks. First, the ASR for PANTS-robustified models doesn’t drop to 0, meaning that these attack methods can still find adversarial samples against the robustified models, albeit less. Second, PANTS assessment is empirical and does not provide theoretical guarantees that the robustified models are robust against the full spectrum of perturbation.
  5 PANTS is practical and efficient to use, generating ~1.7 samples/sec, which is 8X faster than Amoeba.
Refer to caption (a) End-host attacker
Refer to caption (b) In-path attacker
Figure 12: The generation speed for PANTS and Amoeba under different ML models, threat models and applications. PANTS is 8X faster than Ameoba when generating samples.
Thanks to PANTS optimizations for improving the scalability of the SMT solver, including chunking, parallelization, and capability approximations as discussed in §4.4, PANTS is both practical and efficient. Fig. 12 shows the generation speed for PANTS and Amoeba under our two threat models and all applications and implementations. PANTS generates ~1.7 samples/sec, that is 8X faster than Ameoba which generates ~0.2 samples/sec. We acknowledge that this comparison is nuanced, as (unlike PANTS) Amoeba requires training a model prior to generating adversarial samples. To ensure fairness, we report the amortized time, for generating adversarial examples for as many inputs as there are in each dataset and we run experiments in exactly the same machines (described in Appendix in §E). While the overall speed of PANTS clearly outperforms Amoeba, we observe that PANTS’s speed for MLP and RF under the in-path threat model is relatively slow (~0.3 samples/sec). This is because the in-path threat model, together with complex feature engineering, introduces a more complex set of constraints and narrows the range of potential solutions.
  6 Transferability provides an opportunity to robustify a MNC without white-box access.
@@
Refer to caption Figure 13: Transferability results for end-host threat model for the application APP using PANTS’ generated adversarial samples. Most of the ASR are higher than 50%, indicating that the adversarial samples generated for one implementation of an MNC are adversarial for others.
@@
There are use cases in which a network operator will not have access to the MNC, e.g., because the MNC is designed by a third party and its details are secret. While this is not our target setting, thanks to the transferability of adversarial samples PANTS might be able to help the network operator e.g., by providing examples in which the target MNC misclassified.
@@
Transferability is the fraction of adversarial samples, crafted for one implementation of a MNC, that successfully deceive another implementation. Concretely, suppose we evaluate the transferability of the adversarial samples generated from $M_{A}$ to $M_{B}$, we first construct the test set $X_{test\_A\_B}^{\prime}\subseteq X_{test}$ that both $M_{A}$ and $M_{B}$ can correctly classify. Then we use PANTS to generate adversarial samples for $X_{test\_A\_B}^{\prime}$ using $M_{A}$ and evaluate its adversariness against $M_{B}$. Fig. 13 shows the transferability result between MLP, RF, TF and CNN for end-host threat model tested for the application APP. Most of the ASR (outside the diagonal) is above 50%, showing that PANTS can be useful even without full access to the MNC implementation. To benefit from it, the operator would need to train a proxy model and find adversarial samples against it using PANTS. According to Fig. 13, the adversarial samples against the proxy model are likely to be adversarial against the target one.
@@ 6 Conclusion
Acknowledgement
We thank the anonymous reviewers for their informative and insightful feedback. This work was supported by the National Science Foundation (NSF) through Grant CNS-2319442, and a Princeton NextG Innovation Award.
Ethics Considerations
We evaluate PANTS and all baselines on public datasets avoiding any privacy concerns. PANTS and other baselines are not used to perturb live network traffic or an online system, thus avoiding ethical concerns. Our primary goal in proposing PANTS is to assist network operators in evaluating and enhancing the robustness of MNCs to defend against adversaries.
Open Science
We release all artifacts.
First, we release all code and scripts from data processing, MNC training and inference, and PANTS (SMT, AML, including integration with PGD and ZOO) via both Zenodo (https://zenodo.org/records/14728507) and GitHub (https://github.com/jinminhao/PANTS). In terms of secondary artifacts, we share all parameters and configurations used for MNC training and PANTS. We also release the trained MNC models, including both vanilla and robustified versions.
Furthermore, we provide adversarial samples generated by PANTS for each application and model as references. Any artifacts related to the model and data are shared via Zenodo.
We already use public datasets to ensure reproducibility, and the references to these datasets are provided in the paper. That said, we also release the dataset after pre-preprocessing.
Appendix A Robustification via synthetic data augmentation
(a) Accuracy of MLP
(b) Vulnerability of MLP
(c) Accuracy of TF
(d) Vulnerability of TF
Figure 14: Model (MLP and TF) accuracy and ASR when using synthetic data by NetShare and adversarial samples by PANTS for robustification.
Fig. 14 shows the accuracy and vulnerability of MLP and transformer when robustifying using synthetic data generated by NetShare.
Appendix B Features used in different applications
@@
Appendix C Hyperparemeters of ML models
@@
Appendix D Capability extension of Amoeba
@@
Appendix E Testbed specification
For consistent evaluation, all the experiments using PANTS to generate adversarial samples are running on the cluster of Cloudlab machines [19]. Each machine has two Intel Xeon Silver 4114 10-core CPUs at 2.20 GHz and 192GB ECC DDR4-2666 memory. The rest of the experiments related to Amoeba rely on GPU training, which is running on two Azure machines [3]. Each one uses 16 vCPUs from AMD EPYC 7V12(Rome) CPU, 110GB memory, and an Nvidia Tesla T4 GPU with 16 GB GPU memory. Please note that the training and inference of Amoeba is on the CPU machines (same as PANTS) when comparing the computation overhead with PANTS.
Appendix F Robustness of the PANTS- and Amoeba-robustified models
Fig. 15 shows the robustness of using PANTS, Amoeba and BAP to attack PANTS- and Amoeba-robustified models under the in-path threat model.
@@ Appendix F Robustness of the PANTS- and Amoeba-robustified models
Figure 15: ASR for the vanilla, PANTS-robustified and Amoeba-robustified models for the in-path threat model. Similar to the observations in fig. 10, PANTS-robustified’s ASR is lower, demonstrating PANTS’s ability to generate adversarial samples that effectively robustify target models. For the dataset QOE, using Amoeba for robustification is skipped since it cannot effectively generate enough adversarial samples.